OpenStudy (anonymous):

what number can be added to both the numerator and denominator of 61/78 to produce a fraction equal to 4/5?

OpenStudy (anonymous):

7 ?

OpenStudy (anonymous):

\[\frac{61+x}{78+x} = \frac{4}{5} \Rightarrow 5(61+x) = 4(78+x) \Rightarrow 305+5x = 312+4x\] \[x = 7\]

myininaya (myininaya):

i read the question wrong

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

it happens >.< just like my last mishap in the other thread <.<

myininaya (myininaya):

gj you two

OpenStudy (anonymous):

you get medals for trying to make the problem more awesome than it is!

myininaya (myininaya):

lol